Southside Pizzeria wants to improve its ability to manage the ingredient costs associated with making and selling its pizzas. For the month of June, the company plans to make 1,000 pizzas. It has created a planning budget that includes a cost formula for mozzarella cheese of $2.40 per pizza. At the end of June, Southside actually sold 1,100 pizzas and the actual cost of the cheese that it used during the month was $2,632. Required: 1. What is the mozzarella cheese activity variance for June? 2. What is the mozzarella cheese spending variance for June? 3. Assume that the company establishes a price standard of $0.30 per ounce for mozzarella cheese and a quantity standard of eight ounces of cheese per pizza. Also, assume that Southside actually used 9,400 ounces of cheese during the month to make 1100 pizzas. a. What is the materials price variance for mozzarella cheese for June? b. What is the materials quantity variance for mozzarella cheese for June? c. What is the materials spending variance for mozzarella cheese for June?
Answer #1
1
Activity variance $240 U =(1100*2.4)-(1000*2.4)
2
Spending variance $8 F =2632-(1100*2.4)
3
Materials price variance $188 F =2632-(9400*0.3)
Materials Quantity variance $180 U =0.3*(9400-1100*8)
Materials spending variance $8 F =2632-(1100*0.3*8)